A box contains 21 yellow, 28 green and 38 red jelly beans. If 14 jelly beans are selected at random, find each of the following probabilities.
a) The probability that 2 are yellow is: .
b) The probability that 2 are yellow and 11 are green is:
c) The probability that at least one is yellow is:

Answer by jim_thompson5910(35256)

Final answer to each part is in red

For each part, I'm using the formula

n C r = (n!)/(r!*(n-r)!)

which is the combination formula

For example, if n = 10 and r = 2, then

10 C 2 = (10!)/(2!*(10-2)!)

10 C 2 = (10!)/(2!*8!)

10 C 2 = (10*9*8!)/(2!*8!)

10 C 2 = (10*9)/(2!)

10 C 2 = (10*9)/(2*1)

10 C 2 = 45

This problem uses this combination formula and idea extensively. Because of this and because you'll be using this formula a lot in statistics, it will help to know how to compute n C r for any n,r combo (either by hand, table/tree, or a calculator). I won't go into the steps used to compute each combination since there are a lot of them below, but let me know if you need any extra help finding them.



Anyways, onto the actual problem

a)

There are 21 C 2 ways to pick 2 yellow and 66 C 12 ways to pick the remaining colors. This is out of 87 C 14 ways total (to pick 14 jelly beans of any color and any count)

Note: 21+28+38 = 87

So the probability is

((21 C 2)*(66 C 12))/(87 C 14) = 0.19113709882654


-----------------------------------------------------------------------

b)


There are 21 C 2 ways to pick 2 yellow, 28 C 11 ways to pick 11 green, and 38 C 1 ways to pick red. This is out of 87 C 14 ways total

So the probability is

((21 C 2)*(28 C 11)*(38 C 1))/(87 C 14) = 0.00003168297624

-----------------------------------------------------------------------

c)

There are 28+38 = 66 jelly beans that are NOT yellow. So there are 66 C 14 ways to pick 14 jelly beans that are NOT yellow. There are 87 C 14 ways to pick 14 jelly beans (that include yellow).

So the probability of picking 14 jelly beans...and not a single one of them is yellow...is

(66 C 14)/(87 C 14) = 0.014312778044

Now subtract this from one to get

1 - 0.014312778044 = 0.985687221956


So the probability that at least one is yellow is 0.985687221956

Why did I subtract that second to last result from 1? Because the probabilities P(at least one yellow) and P(no yellows) add to 1 (those two events are complementary events...one or the other must happen...but not both at the same time)

So

P(at least one yellow) + P(no yellows) = 1

----> P(at least one yellow) = 1 - P(no yellows)